Grade 12th passPhysical Chemistry

densities of 2 gases are in the ratio 1:2 and their temperatures are in the ratio 2:1. the ratio of their respective pressures is

battleforlife
8 Years agoGrade 12th pass

1 Answer

Arun
8 Years ago
Dear student
 
As we know
P *m = d *R *T
Now
P1/P2 = d1 * T1/ d2 * T2
P1/P2 = (1/2 )* (2/1) = 1/1
Hence
P1/P2 = 1/1
 
Regards
Arun(askIITians forum expert)